OpenStudy (anonymous):

Input the equation of the given line in slope-intercept form. The line through (1, 0) and parallel to x - y = 7.

OpenStudy (anonymous):

well you need a slope. and you are given the equation x-y=7... In the form y=mx+b the m is the slope... So lets solve for y subtract x -y=-x-7 divide by -1 y=x-7... There is not an "m" here so we assume that it is one. So the slope is one and a line parallel to it also has a slope of one .. So now just enter the slope and point in the point slope formula: y-y1=m(x-x1)

OpenStudy (anonymous):

im still confused